This section contains 5 questions. The answer to each of the questions is a single digit integer ranging from 0 to 9.


 






The period of oscillation of a simple pendulum is 2p. Measured value of L is 20.0 cm known to 1 mm accuracy and time for 100 oscillations of the pendulum is found to be 90 s using a wrist watch of 1 s resolution. What is the percentage error in the determination of g?



 

Grade:upto college level

1 Answers

Radhika Batra
10 years ago

3

 

        g = 4p2L/T2;

 

        Here, T =  and DT = . Therefore,

 

        This errors in both L and t are the least count errors. Therefore

 

        (Dg/g) = (DL/L) + 2 (DT/T)

 

        =  = 0.027

 

        Thus, the percentage error in g is

 

        100 (Dg/g) = 100(DL/L) + 2 × 100(DT/T) = 3%